HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

ON-SITE & REAL TIME THz MONITORING OF GASEOUS EMISSION FROM A WASTE RECOVERY CENTER

J. DECKER, F. HINDLE, E. FERTEIN, N. HOUZEL, \underline{G. MOURET} , \underline{A. CUISSET}, Laboratoire de Physico-Chimie de l’Atmosphère, UR4493,

Université du Littoral Côte d’Opale, Dunkerque, France; F. CAZIER, Centre Commun de Mesures, Université du Littoral Côte d’Opale, Dunkerque, France; J.

DUMONT, Paprec Energies 59, Paprec, Dunkerque, France

The Dunkirk waste recovery center continuously treats all of the urban community’s household waste by converting it into electrical and thermal energy through incineration. These activities are carried out with a controlled environmental impact, based on continuous improvement. The TeraWaste project explores the potential of high-resolution TeraHertz spectroscopy for continuous monitoring of gaseous emissions by developing an on-site diagnostic unit. The selectivity and the ability to detect in scattering media make it a relevant alternative to current multi-compounds monitoring solutions. In this context, we have dedicated a sub-millimeter wave source based on a frequency multiplication chain for industrial monitoring applications.^{1} The spectrometer has been characterized by determining the limits of detection of the regulated polar pollutants. Rotational spectroscopy measurements on real process and emission gases, sampled and analyzed off-line as well as in real time on-site, were performed over several hours and compared with standard reference methods and the certified continuous measurement system. Using an adapted preconcentration/thermal desorption setup^{2}^{,}^{3}, a quantified multi-species mapping, extended to the various polar compounds absorbing the sub-mm waves of the plant’s gaseous emissions, in particular VOCs by discriminating them, will enable the operator to better meet current standards and anticipate potential evolution in the regulations.

1doi:10.1109/JSEN.2012.2227055, G. Mouret et al. Versatile Sub-THz Spectrometer for Trace Gas Analysis, IEEE Sens. J., 13(1), 133-138 (2013).

2doi:10.1109/JSEN.2012.2195487, C. F. Neese et al. Compact Submillimeter/Terahertz Gas Sensor With Efficient Gas Collection, Preconcentration, and ppt Sensitivity, IEEE Sens. J., 12(8), 2565-2574 (2012).

3doi:10.3390/s19122719, N. Rothbart et al. Analysis of Human Breath by Millimeter-Wave/Terahertz Spectroscopy, Sensors, 19(12), 2719 (2019).

p-number: p145  Submitted on Wed, 14 Jun 2023 18:49:38 +0200